Information Related to Various Bond Issues) Karen Austin Inc. has issued three types of debt on January 1, 2020, the start of the company's fiscal year.

a.   $10 million, 10-year, 15% unsecured bonds, interest payable quarterly. Bonds were priced to yield 12%.

b.   $25 million par of 10-year, zero-coupon bonds at a price to yield 12% per year.

c.   $20 million, 10-year, 10% mortgage bonds, interest payable annually to yield 12%.

Instructions

Prepare a schedule that identifies the following items for each bond: (1) maturity value, (2) number of interest periods over life of bond, (3) stated rate per each interest period, (4) effective-interest rate per each interest period, (5) payment amount per period and (6) present value of bonds at date of issue.

The question that I want to change: a) revised that the interest is payable quarterly; b and c the calculation for the mortgage payable is the same as for a bond payable
Answer #1
A B C D
1 Unsecured bonds Zero-coupon bonds Mortgage bonds
2 Maturity value 10,000,000 25,000,000 20,000,000
3 Number of interest periods over life of bond 40 10 10
4 Stated rate per each interest period 0.0375 0 10%
5 Effective interest rate per each interest period 0.03 12% 12%
6 Payment amount per period 375000 0 2000000
7 Present value of bonds at date of issue $11,733,607.90 $8,049,330.91 $17,739,910.79

Above figures have been calculated in the following manner:

Unsecured bonds Zero-coupon bonds Mortgage bonds
Maturity value 10000000 25000000 20000000
Number of interest periods over life of bond =10*4 10 10
Stated rate per each interest period =15%/4 0 0.1
Effective interest rate per each interest period =12%/4 0.12 0.12
Payment amount per period =B2*B4 0 =D2*D4
Present value of bonds at date of issue =PV(B5,B3,-B6,0,0)+PV(B5,B3,0,-B2,0) =PV(C5,C3,0,-C2,0) =PV(D5,D3,-D6,0,0)+PV(D5,D3,0,-D2,0)